OpenStudy (anonymous):

Subtract (3x - 30) from the sum of (-2x + 1) and (9x - 9). Choose one answer. a. 10x - 38 b. -4x - 22 c. 8x + 40 d. 4x + 22

Directrix (directrix):

d. 4x + 22

Directrix (directrix):

(-2x + 1) + (9x - 9) - (3x - 30) = 7x -8 - (3x - 30) = 7x - 8 - 3x + 30 = 4x +22